OpenStudy (anonymous):

James drives 5 miles per hour slower than Ken. If James travels 100 miles in the same time that Ken travels 110miles, find each rate. (need an equation to solve it..any help?)

OpenStudy (saifoo.khan):

D=RT 100=(R-5)T OR T=100/(R-5) FOR JAMES 110=RT OR T=110/R FOR KEN SET THESE 2 EQUATIONS EQUAL & SOLVE FOR R. 100/(R-5)=110/R CROSS MULTIPLY, 110(R-5)=100R 110R-550=100R 100R-100R=550 10R=550 R=550/10 R=55 MPH FOR KEN. 55-5=50 MPH IS THE RATE FOR JAMES. 110=55*T T=110/55 T=2 HOURS.